# 3DGraphLLM: Combining Semantic Graphs and Large Language Models for 3D Scene Understanding

### v2

@@ 3.2 Flat Graph Representation
The scene graph is a complete graph because we can generate connections between all pairs of objects. However, such a graph contains $n\cdot(n-1)$ edges between objects, and using the complete graph as a sequence for the LLM would significantly increase the sequence length. However, intuitively, the most relevant relationships for answering user questions are those between an object and its nearest neighbors. Therefore, for each object, we consider a subgraph of its k nearest neighbors. The relationships between objects are encoded using features extracted from point clouds $\{F_{i}^{v}\}_{i=1}^{n}$ and semantic relations features $\{F_{ij}^{e},i\in\{1,...,n\},j\in\{1,...,n\}\}$, represented as a triplet $(F_{i}^{v},F_{ij}^{e},F_{j}^{v})$.
When using the complete scene graph, the number of tokens required to describe the scene is $2\cdot n+3n\cdot(n-1)$. For $100$ objects, which matches the number of objects in the Mask3D (Schult et al., 2023) instance segmentation, this totals $29900$ tokens. By using a $k$-nearest neighbor subgraph, we reduce the token count to $n+3n\cdot k$. As shown in Section 4.3 (see Figure 5), setting $k=2$ improves accuracy in 3D visual-language tasks while reducing the number of tokens needed to describe a scene with $100$ objects to $800$.
Prompt template. Thus, we integrate the scene description as a sequence of object subgraphs into the prompt for LLM in the following way, similar to the integration of the list of objects and their embeddings in the Chat-Scene method (Huang et al., 2024). An example of a prompt for LLM containing a system prompt, a scene description in the form of an object identifier and an object subgraph, a user request, and an LLM assistant response is given in Table 1. The sequence describing an object $i$ starts with its identification token <OBJi>. Then there are $k$ triplets $\{(F_{i}^{v},F_{ij_{k}}^{e},F_{j_{k}}^{v})\}_{j_{k}=1}^{k}$ describing the relationship between the object and its $k$ nearest neighbors.
@@
User: According to the given description, there are brown wooden cabinets,
placed on the side of the kitchen, please provide the ID of the object that closely matches this description.
@@ 3.3 Training Strategy
Following the strategy used in Chat-Scene(Huang et al., 2024), we implement a training approach that involves simultaneously training the projection layers and the language model. We also conduct joint training for various tasks, including visual grounding (ScanRefer (Chen et al., 2020) and Multi3DRefer (Zhang et al., 2023)), 3D scene description (Scan2Cap (Chen et al., 2021)), and 3D visual question answering (ScanQA (Azuma et al., 2022) and SQA3D (Ma et al., 2022)). This adaptation of the tasks is designed for user-assistant interactions, as proposed by the authors of Chat-Scene. During training, we aim to optimize the trainable parameters $\theta$ of both the language model and the projection layers to minimize the negative log-likelihood of the target response $s^{\text{res}}$ compared to the response predicted by the model. We use the loss function from the Chat-Scene method, adapting it to fit our proposed graph representation of the scene given the input prefix sequence $s^{\text{prefix}}$ containing system and user prompts:
@@
where $\ell$ is the length of the token sequence in the LLM response, $s^{\text{res}}_{[1,...,i-1]}$ is the sequence generated up to the $i$-th token. The trainable parameters $\theta$ include the parameters of 3D Object Projection and Semantic Relation Projection Layers, added object identifier token embeddings and the language model.
We use the encoder for semantic relationships between objects pre-trained using ground-truth (GT) point cloud scene segmentation data (Wang et al., 2023a). Since the predicted point cloud segmentation typically contains more noise than the GT segmentation, we anticipate that the edge features derived from the GT segmentation will be of higher quality than those from the neural network instance segmentation. To address this problem, we employ a two-stage training strategy for 3DGraphLLM. First, we pre-train the projection layers and the language model on the GT instance segmentation data to achieve effective projections of the semantic embeddings of relations and objects into the language model’s embedding space. Then, we fine-tune 3DGraphLLM using the noisy data from the neural network segmentation.
@@ 4 Experiments
Datasets. We conduct experiments using the ScanNet (Dai et al., 2017) and 3RScan (Wald et al., 2019) scene datasets. For training 3DGraphLLM on ScanNet scenes (Dai et al., 2017), we utilize data from five 3D vision-language benchmarks: visual grounding tasks (ScanRefer (Chen et al., 2020), Multi3DRefer (Zhang et al., 2023)), scene description (Scan2Cap (Chen et al., 2021)), and 3D visual question answering (ScanQA (Azuma et al., 2022), SQA3D (Ma et al., 2022)). Each of these datasets follows a standard split into training and validation sets, corresponding to 1201 training scans and 312 validation scans from ScanNet. Additionally, we include the RioRefer dataset (Miyanishi et al., 2024), which provides referring expressions for objects in 3RScan scenes (Wald et al., 2019) splitting into standard training and validation sets (1175 training scans and 157 validation scans). Since our method primarily targets visual grounding tasks, the majority of validation experiments are performed on the ScanRefer, Multi3DRefer, and RioRefer datasets.
Implementation details. The projection layers for 3D object features and their semantic relations are three-layer MLPs. In our experiments, we use LLAMA3-8B-Instruct (AI@Meta, 2024), a state-of-the-art large language model, as well as Vicuna-1.5-7B (Zheng et al., 2023) for ablation. For fine-tuning the language model, we apply LoRA (Hu et al., 2021) with a rank of 16. We use a batch size of 8 and train 3DGraphLLM for $3$ epochs with an initial learning rate of 0.00002, following a cosine annealing schedule. Training is performed on a server equipped with an NVIDIA A100 GPU, and the entire training process takes approximately 72 hours. In our experiments, we select $k=2$ nearest neighbors to construct object subgraphs and, in the case of using Mask3D (Schult et al., 2023) instance scene point cloud segmentation, we use a NMS filter and a filter that ensures a minimum distance between nearest neighbors of 1 cm (see Section 4.3).
Evaluation metrics. For the visual grounding task on the ScanRefer (Chen et al., 2020) and RioRefer (Miyanishi et al., 2024) datasets, we use the standard metrics Acc@0.25 and Acc@0.5. A prediction is considered a true positive if the intersection-over-union (IoU) between the predicted object’s 3D bounding box and the ground truth exceeds the thresholds of 0.25 and 0.5, respectively. The Multi3DRefer (Zhang et al., 2023) dataset contains queries that may refer to multiple objects. Therefore, we use the benchmark-standard F1 score at IoU thresholds of 0.25 and 0.5. During ablation experiments, we also assess the quality of object descriptions using the Scan2Cap (Chen et al., 2021) benchmark metrics CIDEr@0.5 and BLEU-4@0.5. For the visual question answering task, we follow the validation strategy from Chat3Dv2, applying CIDEr (Vedantam et al., 2015) and BLEU-4 (Papineni et al., 2002) metrics for ScanQA (Azuma et al., 2022), and exact match accuracy (EM) for SQA3D (Ma et al., 2022).
@@
Table 2: Performance comparison of 3DGraphLLM with state-of-the-art approaches for 3D vision-language tasks. "Expert models" use specialized heads to deal with different 3D vision-language tasks. Our approach falls into the category of "LLM-based models" that consider different tasks as different user queries to a generative model. C denotes the CIDEr metric.
@@
ScanRefer Multi3DRefer ScanQA Sqa3D Scan2Cap
Methods A@0.25↑ A@0.5↑ F1@0.25↑ F1@0.5↑ C↑ B-4↑ EM↑ C@0.5↑ B-4@0.5↑
Expert models ScanRefer (Chen et al., 2020) 37.3 24.3 - - - - - -
MVT (Huang et al., 2022) 40.8 33.3 - - - - - - -
3DVG-Trans (Zhao et al., 2021) 45.9 34.5 - - - - - - -
ViL3DRel (Chen et al., 2022) 47.9 37.7 - - - - - - -
M3DRef-CLIP (Zhang et al., 2023) 51.9 44.7 42.8 38.4 - - - - -
Scan2Cap (Chen et al., 2021) - - - - - - - 35.2 22.4
ScanQA (Azuma et al., 2022) - - - - 64.9 10.1 - - -
Sqa3D (Ma et al., 2022) - - - - - - 47.2 - -
3D-VisTA (Zhu et al., 2023) 50.6 45.8 - - 72.9 13.1 48.5 66.9 34.0
BUTD-DETR (Jain et al., 2022) 52.2 39.8 - - - - - - -
PQ3D (Zhu et al., 2025) - 51.2 - 50.1 87.8 - 47.1 80.3 36.0
LLM-based models ZSVG3D (Yuan et al., 2024) 36.4 32.7 - - - - - -
3D-LLM(Flamingo) (Hong et al., 2023b) 21.2 - - - 59.2 7.2 - - -
3D-LLM(BLIP2-flant5) (Hong et al., 2023b) 30.3 - - - 69.4 12.0 - - -
Chat-3D v2 (Huang et al., 2023) 35.9 30.4 - - 77.1 7.3 - - -
Scene-LLM (Fu et al., 2024) - - - 80.0 12.0 54.2 - -
LL3DA (Chen et al., 2023) - - - - 76.8 13.5 - 65.2 36.8
Grounded 3D-LLM (Chen et al., 2024) 47.9 44.1 45.2 40.6 72.7 13.4 - 70.6 35.5
Chat-Scene (Huang et al., 2024) 55.5 50.2 57.1 52.4 87.7 14.3 54.6 77.1 36.3
3DGraphLLM Vicuna-1.5 (ours) 57.0 51.3 60.1 55.4 87.6 12.1 53.1 81.2 36.3
3DGraphLLM LLAMA3-8B (ours) 60.2 54.6 63.0 58.2 83.1 12.5 55.2 82.9 37.8
Figure 3: Qualitative examples of 3DGraphLLM performance on the ScanRefer dataset. For each query, we provide an RGB image from the ScanNet dataset showing the selected object, along with a visualization of the RGB point cloud. In the point cloud, green points indicate the points that 3DGraphLLM identified as corresponding to the object from the text query, while the green box highlights the ground truth (GT) box for the query.
@@ 4.1 Experimental Results
Comparison with state-of-the-art approaches. As shown in Table 2, our method significantly outperforms baseline approaches that use LLMs on the two ScanNet 3D referred object grounding benchmarks, ScanRefer (Chen et al., 2020) and Multi3DRefer (Zhang et al., 2023), as wall on the Scene Captioning benchmark Scan2Cap (Chen et al., 2021). These results highlight the effectiveness of a learnable graph-based scene representation 3D vision-language tasks. It’s worth noting that the performance of our method is comparable to state-of-the-art specialized models with separate heads for different language tasks, such as 3D-VisTA (Zhu et al., 2023), PQ3D (Zhu et al., 2025) and M3DRef-CLIP (Zhang et al., 2023). Notably, 3DGraphLLM demonstrates a clear advantage over PQ3D (Zhu et al., 2025) and M3DRef-CLIP (Zhang et al., 2023) on the Multi3DRefer dataset.
@@
Qualitative results. Figure 3 shows the qualitative results of 3DGraphLLM on the ScanRefer dataset using Mask3D (Schult et al., 2023) instance scene segmentation. In the left part of the figure, 3DGraphLLM correctly identifies the bed on the right and leverages an additional spatial cue - pants that are lying on the bed. In the right part of the figure, 3DGraphLLM distinguishes the black suitcase next to the refrigerator, despite there being another suitcase farther away from the refrigerator in the scene.
4.2 Ablation Studies. Role of Semantic Relations and Training Pipeline
@@
To isolate the impact of using a scene graph representation, we conduct an experiment with different LLMs and training pipelines using Mask3D (Schult et al., 2023)instance segmentation. We train a version of 3DGraphLLM (3DGraphLLM-0) where the scene is represented as a sequence of object identifiers and features extracted by the 2D Object Encoder and the 3D Object Encoder, following the same training pipeline as 3DGraphLLM (3DGraphLLM-2) with two nearest neighbors. The 3DGraphLLM version with zero nearest neighbors serves as a baseline, equivalent to the Chat-Scene approach, which uses the same LLM as 3DGraphLLM. As shown in Table  3, incorporating a scene graph representation significantly improves the performance of the LLMs across all three 3D Vision-Language tasks: visual grounding, scene description, and question answering. However, the effect is more noticeable for the more modern LLAMA3-8B-Instruct. The pre-training on GT instance segmentation data improves the quality of the 3D Referred Object Grounding for LLAMA3-8B-Instruct and Vicuna-1.5-7B. For LLM Vicuna-1.5-7B, pre-training increases the Scene Captioning quality. For LLAMA3-8B-Instruct, pre-training improves the question answering on the Sqa3D dataset. The most interpretable metrics for the role of semantic edges are the accuracy metrics in the 3D Referred Object Grounding problem, so we keep this pre-training as part of the 3DGraphLLM training pipeline.
Number ScanRefer Multi3DRefer ScanQA Sqa3D Scan2Cap
Methods Pre-train of edges Acc@0.5↑ F1@0.5↑ C↑ B-4↑ EM↑ C@0.5↑ B-4@0.5↑
3DGraphLLM-0 Vicuna1.5 0 50.2 52.4 87.7 14.3 54.6 77.1 36.3
3DGraphLLM-2 Vicuna1.5 2 50.1 52.7 92.2 15.5 54.7 80.4 36.9
3DGraphLLM-2 Vicuna1.5 2 51.3 55.4 87.6 12.1 53.1 81.2 36.3
3DGraphLLM-0 LLAMA3-8B 0 52.0 55.1 84.0 15.8 53.8 80.0 37.5
3DGraphLLM-2 LLLAMA3-8B 2 54.3 57.3 87.4 14.9 54.5 85.6 39.6
3DGraphLLM-2 LLLAMA3-8B 2 54.6 58.2 83.1 12.5 55.2 82.9 37.8
@@
ScanRefer
Methods Instance segmentation Number of edges Minimal distance, cm Acc@0.25↑ Acc@0.5↑
3DGraphLLM-0 GT 0 - $48.9$ $48.9$
3DGraphLLM-2 GT 2 0 $\mathbf{54.4}(+5.6\%)$ $\mathbf{54.4}(+5.6\%)$
3DGraphLLM-0 Mask3D 0 - $46.0$ $34.2$
3DGraphLLM-2 Mask3D 2 0 $47.3(+1.3\%)$ $35.6(+1.4\%)$
3DGraphLLM-2 Mask3D 2 1 $48.0(+2.0\%)$ $36.2(+2.0\%)$
3DGraphLLM-2 Mask3D (+ NMS) 2 1 $\mathbf{48.1}(+2.1\%)$ $\mathbf{36.5}(+2.3\%)$
3DGraphLLM-0 OneFormer3D 0 - $45.4$ $34.5$
3DGraphLLM-2 OneFormer3D 2 0 $47.1(+1.7\%)$ $35.7(+1.2\%)$
3DGraphLLM-2 OneFormer3D (+NMS) 2 1 $\mathbf{47.5}(+2.1\%)$ $\mathbf{36.1}(+1.6\%)$
Figure 4: Comparison of Uni3D object features and VL-SAT semantic edge features for the two nearest neighbors (NNs) based on ground-truth (GT) scene segmentation and Mask3D scene segmentation within the ScanNet training set. Left: Uni3D object features are relatively close for GT point clouds and Mask3D point clouds. Center: using the standard approach for selecting NNs to generate VL-SAT features, the features for pairs of Mask3D point clouds differ significantly from those of GT point clouds. Right: after applying a minimum neighbor distance filter for selecting NNs, the VL-SAT features for object pairs from Mask3D instance segmentation align more closely with those from GT instance segmentation.
4.3 Ablation Studies. 3D Scene Graph Representation
We conduct a series of experiments to explore methods for constructing a scene graph representation from a point cloud. In these experiments, we use a frozen version of LLAMA3-8B-Instruct (AI@Meta, 2024), training only the projection layers. We do not introduce new object tokens into the LLM’s dictionary and follow a three-stage training process, including 3D Object Alignment, 3D Scene Alignment, and Instruction Tuning, as outlined in Chat3D (Wang et al., 2023b).
@@
Quality of instance segmentation. We evaluate how the quality of scene segmentation into objects impacts the performance of 3DGraphLLM. As shown in Table 4, even with noisy neural network segmentation, representing the scene as a graph with semantic relationships is still more effective than using a simple list of objects. We conduct experiments with different object proposal methods, including OneFormer3D (Kolodiazhnyi et al., 2024) and Mask3D (Schult et al., 2023), but found no significant difference between them for our task. Therefore, in subsequent experiments, we use the Mask3D method to maintain consistency with the baseline Chat3Dv2 approach.

### v3

@@ 3.2 Flat Graph Representation
The scene graph is a complete graph since we can generate connections between all pairs of objects. Such a graph contains $n\cdot(n-1)$ edges between objects, and using the complete graph as a sequence for the LLM would significantly increase the sequence length. Intuitively, the most relevant relationships for answering user questions are those between an object and its nearest neighbors. Therefore, for each object, we consider a subgraph of its k nearest neighbors. The relationships between objects are encoded using features extracted from point clouds $\{F_{i}^{v}\}_{i=1}^{n}$ and semantic relations features $\{F_{ij}^{e},i\in\{1,...,n\},j\in\{1,...,n\}\}$, represented as a triplet $(F_{i}^{v},F_{ij}^{e},F_{j}^{v})$.
When using the complete scene graph the number of tokens required to describe the scene is $2\cdot n+3n\cdot(n-1)$. For $100$ objects, which matches the number of object proposals in the Mask3D [46] instance segmentation, this totals $29900$ tokens. By using a $k$-nearest neighbor subgraph, we reduce the token count to $2\cdot n+3n\cdot k$. As shown in  Sec. 4.2 (see  Fig. 4) and Supplementary Materials, setting $k=2$ improves accuracy in 3D visual-language tasks while reducing the number of tokens needed to describe a scene with $100$ objects to $800$. We analyze how the number of objects affects inference speed and GPU memory usage in Supplementary Materials.
Prompt template. We integrate the scene description as a sequence of object subgraphs into the prompt for LLM similar to the integration of the list of object embeddings in the Chat-Scene method [25]. An example of a prompt for LLM containing a system prompt, a scene description in the form of an object identifier, a 2D object feature and an object subgraph, a user request, and an LLM assistant response is given in  Tab. 1. The sequence describing an object $i$ starts with its identification token <OBJi> and 2D object feature $F_{i}^{2d}$. Then there are $k$ triplets $\{(F_{i}^{v},F_{ij_{k}}^{e},F_{j_{k}}^{v})\}_{j_{k}=1}^{k}$ describing the relationship between the object and its $k$ nearest neighbors.
@@
User: According to the given description, there are brown wooden cabinets, placed on the side of the kitchen, please provide the ID of the object that closely matches this description.
@@ 3.3 Training Strategy
Following the strategy used in Chat-Scene[25], we implement a training approach that involves simultaneously training the projection layers and the language model. We also conduct joint training for various tasks, including visual grounding (ScanRefer [5], Multi3DRefer [60], RioRefer [36]), 3D scene description (Scan2Cap [9], Nr3D [1], RioRefer [36]), and 3D visual question answering (ScanQA [3], SQA3D [35], 3RQA [26]). This adaptation of the tasks is designed for user-assistant interactions, as proposed by the authors of Chat-Scene. During training, we aim to optimize the trainable parameters $\theta$ of both the language model and the projection layers to minimize the negative log-likelihood of the target response $s^{\text{res}}$ compared to the response predicted by the model. We use the following loss function:
@@
where $\ell$ is the length of the token sequence in the LLM response, $s^{\text{res}}_{[1,...,i-1]}$ is the sequence generated up to the $i$-th token, $s^{\text{prefix}}$ is the input prefix sequence containing system and user prompts. The trainable parameters $\theta$ include the parameters of 2D/3D Object Projections and Semantic Relation Projection Layers, added object identifier token embeddings, and the language model.
We use the semantic relationships encoder [52] pre-trained using ground-truth (GT) point cloud scene segmentation data. Since the predicted point cloud segmentation typically contains more noise than the GT segmentation, we anticipate that the edge features derived from the GT segmentation will be of higher quality than those from the neural network instance segmentation. To address this problem, we employ a two-stage training strategy for 3DGraphLLM. First, we pre-train the projection layers and the language model on the GT instance segmentation data to achieve effective projections of the semantic embeddings of relations and objects into the language model’s embedding space. Then, we fine-tune 3DGraphLLM using the noisy data from the neural network segmentation.  Sec. 4.2 presents the experimental results, demonstrating the effectiveness of two-stage training and comparing different pre-training datasets.
@@ 4 Experiments
ScanRefer Multi3DRefer Scan2Cap ScanQA Sqa3D
Methods 2D features 3D features LLM A@0.25↑ A@0.5↑ F1@0.25↑ F1@0.5↑ C@0.5↑ B-4@0.5↑ C↑ B-4↑ EM↑
Expert models ScanRefer [5] 37.3 24.3 - - - - - - -
MVT [27] 40.8 33.3 - - - - - - -
3DVG-Trans [61] 45.9 34.5 - - - - - - -
ViL3DRel [6] 47.9 37.7 - - - - - - -
M3DRef-CLIP [60] 51.9 44.7 42.8 38.4 - - - - -
Scan2Cap [9] - - - - 35.2 22.4 - - -
ScanQA [3] - - - - - - 64.9 10.1 -
Sqa3D [35] - - - - - - - - 47.2
3D-VisTA [65] 50.6 45.8 - - 66.9 34.0 72.9 13.1 48.5
BUTD-DETR [28] 52.2 39.8 - - - - - - -
PQ3D [66] - 51.2 - 50.1 80.3 36.0 87.8 - 47.1
LLM-based models ZSVG3D [58] GPT4 36.4 32.7 - - - - - -
3D-LLM [21] Flamingo 21.2 - - - - - 59.2 7.2 -
3D-LLM [21] BLIP2-flant5 30.3 - - - - - 69.4 12.0 -
Chat-3D v2 [24] Vicuna-7B-v0 35.9 30.4 - - - - 77.1 7.3 -
Scene-LLM [15] Llama-2-7B - - - - - - 80.0 12.0 54.2
LEO [26] Vicuna-7B-v1.1 - - - - 72.4 38.2 101.4 13.2 50.0
LL3DA [7] OPT-1.3B - - - - 65.2 36.8 76.8 13.5 -
Grounded 3D-LLM [8] Tiny-Vicuna-1B 47.9 44.1 45.2 40.6 70.6 35.5 72.7 13.4 -
Robin3D [31] Vicuna-7B-v1.5 60.8 55.1 64.9 59.7 87.2 38.4 - - 56.0
GPT4Scene-HD [43] Qwen2-VL-7B 50.9 46.4 53.7 50.0 74.4 37.9 89.9 15.9 57.2
GPT4Scene-HDM [43] Qwen2-VL-7B 62.6 57.0 64.5 59.8 86.3 40.6 96.3 15.5 59.4
Chat-Scene [25] (baseline) Vicuna-7B-v1.5 55.5 50.2 57.1 52.4 77.1 36.3 87.7 14.3 54.6
3DGraphLLM (ours) Vicuna-7B-v1.5 58.6 53.0 61.9 57.3 79.2 34.7 91.2 13.7 55.1
3DGraphLLM (ours) LLAMA3-8B-Instruct 62.4 56.6 64.7 59.9 81.0 36.5 88.8 15.9 55.9
@@
Table 2: Performance comparison of 3DGraphLLM with state-of-the-art approaches for 3D vision-language tasks. ”Expert models” use specialized heads to deal with different 3D vision-language tasks. Our approach falls into the category of ”LLM-based models” that consider different tasks as different user queries to a generative model. C denotes the CIDEr metric.
Datasets. For pretraining 3DGraphLLM using GT instance segmentation, we employ a combined 3D Vision-Language dataset for ScanNet [11] and 3RScan [50] scenes. For ScanNet scenes, we utilize data from five 3D vision-language benchmarks: visual grounding tasks (ScanRefer [5], Multi3DRefer [60]), scene description (Scan2Cap [9]), and 3D visual question answering (ScanQA [3], SQA3D [35]). Each of these datasets follows a standard split into training and validation sets, corresponding to 1201 training scans and 312 validation scans from ScanNet. For 3RScan scenes, we use data from the RioRefer dataset [36] for object grounding, and the 3RQA dataset [26] for question answering. For 3RScan data, we follow the standard train/validation scan split and use the scans present in the RioRefer dataset for training, resulting in 1175 training scans and 157 validation scans. To augment the data for the scene description task, we use data from the RioRefer [36] and Nr3D [1] datasets, taking object grounding queries provided in these datasets as reference descriptions of objects in the scene. To assess 3DGraphLLM performance under realistic conditions, we perform fine-tuning on predicted instance segmentation using 3D vision-language benchmarks for ScanNet scenes: ScanRefer, Multi3DRefer, Scan2Cap, ScanQA, and SQA3D.
Implementation details. The projection layers for 2D/3D object features and their semantic relations are three-layer MLPs. In our experiments, we use LLAMA3-8B-Instruct [2], a state-of-the-art large language model, as well as Vicuna-1.5-7B [62] for ablation. For fine-tuning the language model, we apply LoRA [23] with a rank of 16. We use a batch size of 8 and train 3DGraphLLM for $3$ epochs with an initial learning rate of $5\cdot 10^{-6}$, following a cosine annealing schedule. Training is performed on a server equipped with $4$ NVIDIA A100 GPUs, and the entire training process takes approximately $24$ hours. In our experiments, we select $k=2$ nearest neighbors to construct object subgraphs and, in the case of using Mask3D [46] instance scene point cloud segmentation, we use a NMS filter and a filter that ensures a minimum distance between nearest neighbors of 1 cm (see  Sec. 4.2).
Method
Dataset 3DGraphLLM GPT4Scene
Input token number per scene 800 10400
Inference speed, sec ScanRefer 0.4 1.9
Multi3DRefer 0.5 2.0
Scan2Cap 0.9 2.2
ScanQA 0.4 1.9
SQA3D 0.4 1.7
Table 3: Input tokens and inference speed comparison (Mask3D instance segmentation).
Evaluation metrics. For the visual grounding task on the ScanRefer [5]dataset, we use the standard metrics Acc@0.25 and Acc@0.5. A prediction is considered a true positive if the intersection-over-union (IoU) between the predicted object’s 3D bounding box and the ground truth exceeds the thresholds of 0.25 and 0.5, respectively. The Multi3DRefer [60] dataset contains queries that may refer to multiple objects. Therefore, we use the benchmark-standard F1 score at IoU thresholds of 0.25 and 0.5. We assess the quality of object descriptions using the Scan2Cap [9] benchmark metrics CIDEr@0.5 and BLEU-4@0.5. For the visual question answering task, we follow the validation strategy from Chat-Scene[25], applying CIDEr [49] and BLEU-4 [39] metrics for ScanQA [3], and exact match accuracy (EM) for SQA3D [35].
Figure 3: Qualitative examples of 3DGraphLLM performance on object grounding, dense captioning, and question answering tasks. We provide a visualization of the RGB point cloud along with blue objects bounding boxes.
@@
Number Training scenes ScanRefer Multi3DRefer Scan2Cap ScanQA Sqa3D
Methods LLM Pre-train of edges Acc@0.5↑ F1@0.5↑ C@0.5↑ B-4@0.5↑ C↑ B-4↑ EM↑
3DGraphLLM-0 Vicuna1.5-7B 0 ScanNet 50.2 52.4 77.1 36.3 87.7 14.3 54.6
3DGraphLLM-2 Vicuna1.5-7B 2 ScanNet 50.1 52.7 80.4 36.9 92.2 15.5 54.7
3DGraphLLM-2 Vicuna1.5-7B 2 ScanNet+3RScan 53.1 57.3 79.2 34.7 91.2 13.7 55.1
3DGraphLLM-0 LLAMA3-8B-Instruct 0 ScanNet 52.0 55.1 80.0 37.5 84.0 15.8 53.8
3DGraphLLM-2 LLAMA3-8B-Instruct 2 ScanNet 54.3 57.3 85.6 39.6 87.4 14.9 54.5
3DGraphLLM-2 LLAMA3-8B-Instruct 2 ScanNet 56.2 58.7 82.9 37.3 85.4 15.1 55.6
3DGraphLLM-2 LLAMA3-8B-Instruct 2 ScanNet+3RScan 56.6 59.9 81.0 36.5 88.8 15.9 55.9
@@ 4.1 Experimental Results
Comparison with state-of-the-art approaches. As shown in  Tab. 2, our method significantly outperforms the baseline approach Chat-Scene [25] on the two ScanNet 3D referred object grounding benchmarks, ScanRefer [5] and Multi3DRefer [60], as well as on the scene captioning benchmark Scan2Cap [9] and the question answering benchmarks ScanQA [3] and SQA3D [35]. These results highlight the effectiveness of a learnable graph-based scene representation 3D vision-language tasks. It’s worth noting that the performance of our method surpasses state-of-the-art specialized models with separate heads for different language tasks, such as 3D-VisTA [65], PQ3D [66], and M3DRef-CLIP [60].
Notably, 3DGraphLLM demonstrates state-of-the-art quality for the 3D referred object grounding task for LLM-based methods. In particular, our 3DGraphLLM with LLAMA3-8B as the base LLM outperforms Robin3D [31] on ScanRefer benchmark showing comparable quality on Multi3DRefer and SQA3D benchmarks. Robin3D is trained on 1M instruction-following data that are not publicly available, while our approach uses only 370K instruction-following data. Our experiments in  Tab. 4 highlight the importance of training data for 3DGraphLLM, suggesting that incorporating more data for fine-tuning could further improve its performance. 3DGraphLLM achieves results comparable to the state-of-the-art method GPT4Scene-HDM [43], showing the importance of semantic relations for this task. At the same time, 3DGraphLLM uses fewer tokens to describe the scene (see Tab. 3), allowing up to five times faster inference for object-grounding tasks.
@@
Qualitative results.  Fig. 3 shows the qualitative results of 3DGraphLLM using Mask3D [46] instance scene segmentation. 3DGraphLLM efficiently uses spatial cues for solving 3D Vision-Language tasks. For example, 3DGraphLLM distinguishes the black suitcase next to the refrigerator, despite there being another suitcase farther away from the refrigerator in the scene. In Supplementary Materials we provide more examples of 3DGraphLLM performance.
4.2 Ablation Studies
@@
Role of Semantic Relations. To isolate the impact of using a scene graph representation, we conduct an experiment with different LLMs and training pipelines using Mask3D [46] instance segmentation. We train a version of 3DGraphLLM (3DGraphLLM-0) where the scene is represented as a sequence of object identifiers and features extracted by the 2D Object Encoder and the 3D Object Encoder, following the same training pipeline as 3DGraphLLM (3DGraphLLM-2) with two nearest neighbors. The 3DGraphLLM version with zero nearest neighbors serves as a baseline, equivalent to the Chat-Scene approach, which uses the same LLM as 3DGraphLLM-2. As shown in  Tab. 4, incorporating a scene graph representation significantly improves the performance of the LLMs across all three 3D Vision-Language tasks: visual grounding, scene description, and question answering. However, the effect is more noticeable for the more recent LLAMA3-8B-Instruct.
Training pipeline. The pre-training on GT instance segmentation data improves the quality of the 3D referred object grounding for LLAMA3-8B-Instruct and Vicuna-1.5-7B. For LLM Vicuna-1.5-7B, pre-training increases the scene captioning quality. For LLAMA3-8B-Instruct, pre-training improves the question answering on the SQA3D dataset. We compare two pre-training datasets for 3DGraphLLM using LLAMA3-8B-Instruct. The first contains only 3D Vision-Language data from ScanNet, while the second includes data from both ScanNet and 3RScan.  Tab. 4 shows that incorporating 3RScan data further enhances object grounding and question answering performance. The most interpretable metrics for the role of semantic edges are the accuracy metrics in the 3D referred object grounding task, so we keep this pre-training as part of the 3DGraphLLM training pipeline.
It is worth noting that the n-gram-based evaluation metrics used in scene captioning and question answering benchmarks are not adequate for assessing the quality of LLM-generated responses because they fail to capture the flexibility and richness of LLM outputs. This effect is particularly noticeable in the scene captioning task, where CIDEr@0.5 and BLEU-4@0.5 penalize 3DGraphLLM if the model incorporates visual and spatial cues that are missing from the reference descriptions. For example, in the scene shown in  Fig. 3, 3DGraphLLM describes a toilet as: ”This is a white toilet. It is to the right of the shower curtain.” This is a correct description of the object, yet the reference captions use different wording and spatial cues, causing CIDEr@0.5 to assign a score of 0.0 to this description. See Supplementary Materials for a more detailed illustration of this effect.
Instance Number Minimal ScanRefer Multi3DRef
Methods segmentation of edges distance, cm Acc@0.5↑ F1@0.5↑
3DGraphLLM-0 GT 0 - 61.5 64.4
3DGraphLLM-2 GT 2 0 66.9 69.9
3DGraphLLM-0 Mask3D 0 - 52.0 55.1
3DGraphLLM-2 Mask3D 2 0 55.6 58.2
3DGraphLLM-2 Mask3D (+ NMS) 2 0 55.7 58.6
3DGraphLLM-2 Mask3D (+ NMS) 2 1 56.2 58.7
3DGraphLLM-0 OneFormer3D 0 - 50.0 52.8
3DGraphLLM-2 OneFormer3D 2 0 52.8 55.8
3DGraphLLM-2 OneFormer3D (+NMS) 2 1 54.6 57.2
@@
Instance Relations Number ScanRefer Multi3DRef
Methods segmentation as triplets of edges Acc@0.5↑ F1@0.5↑
3DGraphLLM-0 Mask3D 0 52.0 55.1
3DGraphLLM-2 Mask3D 2 54.2 56.3
3DGraphLLM-2 Mask3D 2 54.3 57.3
Table 6: Ablation study on subgraph representation.
@@
Quality of instance segmentation. We evaluate how the quality of scene segmentation into objects impacts the performance of 3DGraphLLM. For these experiments, we use the full training pipeline with a pre-training phase on GT instance segmentation on ScanNet data. As shown in  Tab. 5, even with noisy neural network segmentation, representing the scene as a graph with semantic relationships is still more effective than using a simple list of objects. We conduct experiments with different object proposal methods, including OneFormer3D [33] and Mask3D [46], but we found that Mask3D segmentation shows better results for our tasks. Therefore, in subsequent experiments, we use the Mask3D method to maintain consistency with the baseline Chat-Scene approach.